# The Hallucination Dilemma: Factuality-Aware Reinforcement Learning for Large Reasoning Models

### v1

@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable breakthroughs in natural language generation and complex reasoning tasks [9, 15]. This progress is fueled by new training paradigms that imbue LLMs with reasoning capabilities, including fine-tuning on long chain-of-thought (CoT) solutions [43] and reinforcement learning (RL) based optimization [9]. By leveraging these approaches, state-of-the-art models can decompose complex problems into reasoning steps and iteratively refine their answers, yielding impressive results on challenging benchmarks in mathematics [10], multi-hop question answering [41], coding [2], and decision-making [32].
Along with improved reasoning ability, a critical question has emerged: To what extent do reasoning models maintain factuality after RL fine-tuning? In this work, our empirical observations indicate that when large LLMs are trained with outcome-driven RL, they tend to produce more factually incorrect or fabricated statements (i.e., hallucinations [13, 46]) during their reasoning process. We find that these hallucinations usually stem from incorrect intermediate reasoning, i.e., the model may generate unsupported or false claims in the chain-of-thought, even if its final answer is occasionally correct. Indeed, RL-aligned models can be highly fluent and confident in their step-by-step explanations while subtly introducing factual errors [40]. This issue undermines the reliability of RL-trained reasoning models, as a flawed or untruthful answer is problematic for trust and interpretability.
@@ 2 Preliminary Experiments
Models and Datasets. For comparative analysis, we select several groups of LLMs without and with large-scale RL training: (1) DeepSeek-V3 [21] and DeepSeek-R1 [9]; (2) Qwen2.5-32B-Instruct [39] and QwQ-32B [28]; (3) Qwen2.5-14B/7B/1.5B-Instruct [39] and R1-Distill-Qwen-14B/7B/1.5B [9]; (4) Llama3.1-8B-Instruct [8] and R1-Distill-Llama-8B [9]. These distilled models are fine-tuned on the synthesized long-CoT data from DeepSeek-R1, thus exhibiting similar thinking modes to RL-tuned models. For the hallucination datasets, we select three benchmarks, namely TruthfulQA [20], HaluEval [19], and HalluQA [4]. In TruthfulQA, we adopt the generation task, where the model is prompted to generate an answer to a question, and then utilize a fine-tuned GPT-3 to predict the truthfulness of the answer. The hallucination rate is equal to the ratio of untruthful answers. HalluQA follows the same method as TruthfulQA, except that it employs GPT-4 as the evaluator. For HaluEval, we adopt the QA subset, where the LLM is provided with a question and two candidate answers and it needs to assess which answer is factually correct. The hallucination rate refers to the ratio of false judgments made by the model.
@@ 3 Background
Reinforcement Learning. In reinforcement learning, language model generation can be formulated as a token-level Markov decision process, denoted by =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\pi_{\theta}\}$. At the $t$-th generation step, the state $s_{t}\in\mathcal{S}$ is the concatenation of the input prompt $x$ and the generated sequence so far $y_{<t}=[o_{1},...,o_{t-1}]$. The policy $\pi_{\theta}(\cdot|s_{t})$ (parameterized by $\theta$) selects the next action $o_{t}\in\mathcal{A}$ by generating a token from the vocabulary $\mathcal{A}$, leading to a deterministic transition to the next state $s_{t+1}=[s_{t},o_{t}]$. Let $Y^{*}$ denote the set of correct answers for a given input prompt $x$. Existing research mostly adopts a binary reward $\mathcal{R}$ for each output $y$, which is sparse and discontinuous [32, 18]. For simplicity, we consider a prevalent format in this paper: $\mathcal{R}(y)=1$ when $y\in Y^{*}$, or $\mathcal{R}(y)=0$ when $y\notin Y^{*}$. The expected return (reward) under the policy $\pi_{\theta}$ is:
@@
The policy gradient methods maximize the expected total reward by repeatedly estimating the gradient. Several policy gradient algorithms have been proposed such as Proximal Policy Optimization [29].
Group Relative Policy Optimization. In DeepSeek-Math [30], researchers proposed a new policy gradient method, Group Relative Policy Optimization (GRPO), which improves the effectiveness and efficiency of RL training based on Proximal Policy Optimization (PPO) [29]. The PPO algorithm maximizes the following surrogate objective:
@@ 4 Theoretical Analysis
The REINFORCE gradient for a given input is $g=R(y)\,\nabla_{\theta}\log\pi_{\theta}(y|x)$. This is a random vector depending on the sampled output $y$. There is a $p$ chance that the sampled output is correct (yielding $R=1$) and a $1-p$ chance that it is wrong (yielding $R=0$). Hence, the gradient $g$ is zero with probability $1-p$, and equals $\nabla_{\theta}\log\pi_{\theta}(y^{*}|x)$ with probability $p$. The expected gradient is $\mathbb{E}[g]=p\,\nabla_{\theta}\log\pi_{\theta}(y^{*}|x)$. The second moment is $\mathbb{E}[\|g\|^{2}]=p\,\|\nabla_{\theta}\log\pi_{\theta}(y^{*}|x)\|^{2}$. Therefore, the covariance (and in particular the trace as a measure of variance magnitude) is on the order of
@@
For very small $p$, the factor $1-p$ is near 1, so roughly $\text{Var}(g)\approx p\,\|\nabla\log\pi_{\theta}(y^{*})\|^{2}$. In other words, when the model has low chance to generate correct answers, the gradient estimator is almost always zero but occasionally a large spike, yielding extremely high variance relative to its mean. This high variance makes the training updates very unstable, as gradient estimates will fluctuate widely between zero and large values. We conclude that, with a binary reward, most of the time the model receives no learning signal, and when it does, the signal is an outsized jump. ∎
Theorem 4.2 (Exploration–Exploitation Trade-off and Hallucination Risk).
@@
The necessity for exploration to ensure non-stagnation at zero-reward local minima introduces the exploration-exploitation trade-off. On one hand, the entropy must remain relatively high to explore potential answers, increasing the probability of hallucinated outputs. On the other hand, maximizing immediate reward requires concentrating distribution on already known rewarding answers.
@@ 5.1 Reward Function Design
Step-wise Factuality Reward. Typically, factuality is mainly related to world knowledge, so we focus on complex knowledge-intensive tasks in this work, e.g., multi-hop question answering [41]. Specifically, for an input question $x$, we assume that it can be associated with knowledge snippets in a source such as Wikipedia, denoted by $\mathcal{K}$, which can be called evidence. We first split the reasoning text into multiple sentences as steps, denoted by $\{z_{1},...,z_{N}\}$, and then employ an LLM to determine whether the evidence $\mathcal{K}$ can entail each sentence $z_{j}$ to assign a step-wise factuality reward as:
$\mathcal{R}_{\text{factuality}}(z_{j})=\begin{cases}1,&\text{if the sentence $z_{j}$ can be entailed from the evidence $\mathcal{K}$}\\ 0,&\text{if the sentence $z_{j}$ is neutral to the evidence $\mathcal{K}$}\\ -1,&\text{if the sentence $z_{j}$ contradicts the evidence $\mathcal{K}$}\end{cases}$ (8)
@@
Answer Correctness Reward. In most existing research [9, 38, 33], answer correctness reward is usually adopted due to its significant scalability and simplicity. Specially, given an input prompt $x$, the policy model generates an output $y$, which includes intermediate reasoning texts followed by the final answer. We use a rule-based reward function to evaluate an answer’s correctness as follows:
$\mathcal{R}_{\text{answer}}(y)=\begin{cases}1,&\text{if the final answer fullymatches the ground truth}\\ 0,&\text{if the final answer does not match the ground truth}\end{cases}$ (9)
@@ 5.2 Factuality-Aware Policy Optimization
For online policy optimization, we adopt Group Relative Policy Optimization (GRPO) [30], which samples a group of outputs $\{y_{1},...,y_{G}\}$ for each input $x$ and then obtains their rewards $\{\mathcal{R}_{1},...,\mathcal{R}_{G}\}$ to compute the advantage $A_{i}$ for $y_{i}$ as $A_{i}={\left(\mathcal{R}_{i}-\text{mean}(\{\mathcal{R}_{1},...,\mathcal{R}_{G}\})\right)}/{\text{std}(\{\mathcal{R}_{1},...,\mathcal{R}_{G}\})}$.
Factuality-Aware Advantage Adjustment. Most existing work assigns the same advantage (derived from the overall response) to all tokens [30, 38], which is suboptimal and does not differentiate between them. We argue that, regardless of whether the final answer is correct, factually correct reasoning tokens should be encouraged, while incorrect reasoning tokens should be penalized. Therefore, we adjust the advantage of each token $o_{i,t}\in z_{j}$ for the output $y_{i}$ as follows:
$\hat{A}_{i,t}=\begin{cases}A_{i},&\text{if $A_{i}>0\land\mathcal{R}_{\text{factuality}}(z_{j})=1$ or $A_{i}<0\land\mathcal{R}_{\text{factuality}}(z_{j})=-1$}\\ -A_{i},&\text{if $A_{i}>0\land\mathcal{R}_{\text{factuality}}(z_{j})=-1$ or $A_{i}<0\land\mathcal{R}_{\text{factuality}}(z_{j})=1$}\\ A_{i},&\text{if $A_{i}=0$ or $\mathcal{R}_{\text{factuality}}(z_{j})=0$}\\ \end{cases}$ (11)
@@ 6.1 Experimental Setup
Datasets and Metrics. To perform FSPO, we adopt a challenging subset of HotpotQA [41] and 2WikiMultiHopQA [11] from previous work [33] and randomly select 2K examples for training (including questions, answers, and corresponding Wikipedia segments). Besides, to improve the complex reasoning capabilities of LLMs, we incorporate the SimpleRL dataset [45] with 8K mathematical problems for vanilla RL training. For hallucination evaluation, we adopt three widely-used benchmarks, i.e., TruthfulQA [20], HaluEval [19], and HalluQA [4]. For HaluEval, we adopt the QA subset and report the accuracy with which the model can discriminate correct and hallucinated answers. For TruthfulQA and HalluQA, we report the ratio of outputs that the judge model predicts as truthful. We employ GSM8K [5], MATH-500 [10], AIME 2024 [22], and AIME 2025 [23] to evaluate mathematical reasoning performance and report the Pass@1 accuracy results in these datasets. For more details, refer to Appendix C.
Baselines. We employ Qwen2.5-7B-Base/Instruct [39] and Llama3.1-8B-Instruct [8] as backbone models for our approach. We compare our model against three groups of baselines:
$\bullet$ API-based Models include DeepSeek-V3 [21], DeepSeek-R1 [9], OpenAI GPT-4o, and GPT-o1.
$\bullet$ Reasoning Models consist of QwQ-32B [28], Distill-Qwen-7B/14B/32B, and Distill-Llama-8B [9], which are fine-tuned on the long CoT data (800K samples) synthesized from DeepSeek-R1.
@@
Implementation Details. We train our model using the verl [31] framework. Specifically, during training, we use a batch size of 8, generate 8 rollouts for each prompt, set a maximum prompt and response length of 2,048 tokens, and train using a mini-batch size of 1,024. We train our model on the mixture dataset of HotpotQA subset and SimpleRL for one epoch with a constant learning rate of 4e-7 and temperature parameter of 1.0. The coefficient for KL loss is set to 1e-3 and the clip ratio is 0.2. We adopt the HHEM-2.1 model [14] to verify step-wise factuality. During evaluation, we set the sampling temperature to 1.0. We present prompt templates throughout this work in Appendix F.
@@ 6.2 Main Results
The experimental results presented in Table 1 demonstrate the performance of our proposed method, FSPO, across various reasoning and hallucination benchmarks compared to multiple competitive models, including API-based, reasoning, and open-source alternatives. On hallucination benchmarks, FSPO clearly outperforms all open-source, reasoning and even some API-based models. Specifically, FSPO based on Qwen2.5-7B-Base achieves the highest scores among all open-source and reasoning models on hallucination datasets, clearly demonstrating the effectiveness of our training approach in substantially reducing hallucinations and enhancing factual consistency.
@@ 6.3 Further Analysis
RL Algorithm Analysis. We further adapt FSPO to the Reinforce++ algorithm [12] for validating the adaptability and effectiveness of FSPO beyond GRPO. As shown in Figure 5, FSPO with Reinforce++ exhibits significant improvement over Reinforce++ alone, closely following the performance of FSPO with GRPO. Moreover, FSPO with GRPO consistently achieves higher accuracy than the Reinforce++ baselines across training steps on both the MATH-500 and HaluEval-QA benchmarks. These findings indicate that our proposed FSPO approach generalizes well to other reinforcement learning algorithms, enhancing their effectiveness and stability in solving complex reasoning and reducing hallucinations.
@@
Number of Training Samples. To investigate the model performance under varying numbers of training samples, we randomly sample 1K, 2K, 4K, and 8K samples from the HotpotQA subset [33], with the results shown in Figure 7. On MATH-500, the accuracies are similar from 1K to 4K samples while diminishing with 8K samples, which might be because more factual QA data lower the math reasoning performance. In contrast, on the HaluEval-QA task, significant accuracy improvement continues up to 2K samples, with 4K and 8K samples achieving mediocre results. This demonstrates that our approach can significantly reduce hallucinations using only a small amount of data, without compromising the model’s capacity for complex reasoning.

### v2

# Reasoning Models Hallucinate More: Factuality-Aware Reinforcement Learning for Large Reasoning Models

@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable breakthroughs in natural language generation and complex reasoning tasks [deepseek-r1, openai-o1]. This progress is fueled by new training paradigms that imbue LLMs with reasoning capabilities, including fine-tuning on long chain-of-thought (CoT) solutions [limo] and reinforcement learning (RL) based optimization [deepseek-r1]. By leveraging these approaches, state-of-the-art models can decompose complex problems into reasoning steps and iteratively refine their answers, yielding impressive results on challenging benchmarks in mathematics [hendrycksmath2021], multi-hop question answering [hotpotqa], coding [humaneval], and decision-making [alphago].
Along with improved reasoning ability, a critical question has emerged: To what extent do reasoning models maintain factuality after RL fine-tuning? In this work, our empirical observations indicate that when LLMs are trained with outcome-driven RL, they tend to produce more factually incorrect or fabricated statements (i.e., hallucinations [hallucination_survey, Zhang-CoRR-2023-Siren]) during their reasoning process. We find that these hallucinations usually stem from incorrect intermediate reasoning, i.e., the model may generate unsupported or false claims in the chain-of-thought, even if its final answer is occasionally correct. Indeed, RL-aligned models can be highly fluent and confident in their step-by-step explanations while subtly introducing factual errors [yang2025kongzi]. This issue undermines the reliability of RL-trained reasoning models, as a flawed or untruthful answer is problematic for trust and interpretability.
@@ 2 Preliminary Experiments
Models and Datasets. For comparative analysis, we select several groups of LLMs without and with large-scale RL training: (1) DeepSeek-V3 [deepseek-v3] and DeepSeek-R1 [deepseek-r1]; (2) Qwen2.5-32B-Instruct [yang2024qwen2] and QwQ-32B [qwq32b]; (3) Qwen2.5-14B/7B/1.5B-Instruct [yang2024qwen2] and R1-Distill-Qwen-14B/7B/1.5B [deepseek-r1]; (4) Llama3.1-8B-Instruct [grattafiori2024llama] and R1-Distill-Llama-8B [deepseek-r1]. These distilled models are fine-tuned on the synthesized long-CoT data from DeepSeek-R1, thus exhibiting similar thinking modes to RL-tuned models. For the hallucination datasets, we select three benchmarks, namely TruthfulQA [truthfulqa], HaluEval [halueval], and HalluQA [halluqa]. In TruthfulQA, we adopt the generation task, where the model is prompted to generate an answer to a question, and then utilize a fine-tuned GPT-3 to predict the truthfulness of the answer. The hallucination rate is equal to the ratio of untruthful answers. HalluQA follows the same method as TruthfulQA, except that it employs GPT-4 as the evaluator. For HaluEval, we adopt the QA subset, where the LLM is provided with a question and two candidate answers and it needs to assess which answer is factually correct. The hallucination rate refers to the ratio of false judgments made by the model.
@@ 3 Background
Reinforcement Learning. In reinforcement learning, language model generation can be formulated as a token-level Markov decision process, denoted by $\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{R},\pi_{\theta}\}$. At the $t$-th generation step, the state $s_{t}\in\mathcal{S}$ is the concatenation of the input prompt $x$ and the generated sequence so far $y_{<t}=[o_{1},...,o_{t-1}]$. The policy $\pi_{\theta}(\cdot|s_{t})$ (parameterized by $\theta$) selects the next action $o_{t}\in\mathcal{A}$ by generating a token from the vocabulary $\mathcal{A}$, leading to a deterministic transition to the next state $s_{t+1}=[s_{t},o_{t}]$. Let $Y^{*}$ denote the set of correct answers for a given input prompt $x$. Existing research mostly adopts a binary reward $\mathcal{R}$ for each output $y$, which is sparse and discontinuous [alphago, ladosz2022exploration]. For simplicity, we consider a prevalent format in this paper: $\mathcal{R}(y)=1$ when $y\in Y^{*}$, or $\mathcal{R}(y)=0$ when $y\notin Y^{*}$. The expected return (reward) under the policy $\pi_{\theta}$ is:
@@
The policy gradient methods maximize the expected total reward by repeatedly estimating the gradient. Several policy gradient algorithms have been proposed such as Proximal Policy Optimization [schulman2017proximal].
Group Relative Policy Optimization. In DeepSeek-Math [deepseek-math], researchers proposed a new policy gradient method, Group Relative Policy Optimization (GRPO), which improves the effectiveness and efficiency of RL training based on Proximal Policy Optimization (PPO) [schulman2017proximal]. The PPO algorithm maximizes the following surrogate objective:
@@ 4 Theoretical Analysis
The REINFORCE gradient for a given input is $g=R(y)\,\nabla_{\theta}\log\pi_{\theta}(y|x)$. This is a random vector depending on the sampled output $y$. There is a $p$ chance that the sampled output is correct (yielding $R=1$) and a 1p1-p chance that it is wrong (yielding $R=0$). Hence, the gradient $g$ is zero with probability 1p1-p, and equals $\nabla_{\theta}\log\pi_{\theta}(y^{*}|x)$ with probability $p$. The expected gradient is $\mathbb{E}[g]=p\,\nabla_{\theta}\log\pi_{\theta}(y^{*}|x)$. The second moment is $\mathbb{E}[\|g\|^{2}]=p\,\|\nabla_{\theta}\log\pi_{\theta}(y^{*}|x)\|^{2}$. Therefore, the covariance (and in particular the trace as a measure of variance magnitude) is on the order of
@@
For very small $p$, the factor 1p1-p is near 1, so roughly $\text{Var}(g)\approx p\,\|\nabla\log\pi_{\theta}(y^{*})\|^{2}$. In other words, when the model has low chance to generate correct answers, the gradient estimator is almost always zero but occasionally a large spike, yielding extremely high variance relative to its mean. This high variance makes the training updates very unstable, as gradient estimates will fluctuate widely between zero and large values. We conclude that, with a binary reward, most of the time the model receives no learning signal, and when it does, the signal is an outsized jump. ∎
Theorem 4.2 (Exploration-Exploitation Trade-off and Hallucination Risk).
@@
The necessity for exploration to ensure non-stagnation at zero-reward local minima introduces the exploration-exploitation trade-off. On one hand, the entropy must remain relatively high to explore potential answers, increasing the probability of hallucinated outputs. On the other hand, maximizing immediate reward requires concentrating distribution on already known rewarding answers. ∎
@@ 5.1 Reward Function Design
Step-wise Factuality Reward. Typically, factuality is mainly related to world knowledge, so we focus on complex knowledge-intensive tasks in this work, e.g., multi-hop question answering [hotpotqa]. Specifically, for an input question $x$, we assume that it can be associated with knowledge snippets in a source such as Wikipedia, denoted by $\mathcal{K}$, which can be called evidence. We first split the reasoning text into multiple sentences as steps, denoted by $\{z_{1},...,z_{N}\}$, and then employ an LLM to determine whether the evidence $\mathcal{K}$ can entail each sentence $z_{j}$ to assign a step-wise factuality reward as:
[…]
@@
Answer Correctness Reward. In most existing research [deepseek-r1, xie2025logic, r1-searcher], answer correctness reward is usually adopted due to its significant scalability and simplicity. Specially, given an input prompt $x$, the policy model generates an output $y$, which includes intermediate reasoning texts followed by the final answer. We use a rule-based reward function to evaluate an answer’s correctness as follows:
[…]
@@ 5.2 Factuality-Aware Policy Optimization
For online policy optimization, we adopt Group Relative Policy Optimization (GRPO) [deepseek-math], which samples a group of outputs $\{y_{1},...,y_{G}\}$ for each input $x$ and then obtains their rewards $\{\mathcal{R}_{1},...,\mathcal{R}_{G}\}$ to compute the advantage $A_{i}$ for $y_{i}$ as $A_{i}={\left(\mathcal{R}_{i}-\text{mean}(\{\mathcal{R}_{1},...,\mathcal{R}_{G}\})\right)}/{\text{std}(\{\mathcal{R}_{1},...,\mathcal{R}_{G}\})}$.
Factuality-Aware Advantage Adjustment. Most existing work assigns the same advantage (derived from the overall response) to all tokens [deepseek-math, xie2025logic], which is suboptimal and does not differentiate between them. We argue that, regardless of whether the final answer is correct, factually correct reasoning tokens should be encouraged, while incorrect reasoning tokens should be penalized. Therefore, we adjust the advantage of each token $o_{i,t}\in z_{j}$ for the output $y_{i}$ as follows:
[…]
@@ 6.1 Experimental Setup
Datasets and Metrics. To perform FSPO, we adopt a challenging subset of HotpotQA [hotpotqa] and 2WikiMultiHopQA [2wikimultihopqa] from previous work [r1-searcher] and randomly select 2K examples for training (including questions, answers, and corresponding Wikipedia segments). Besides, to improve the complex reasoning capabilities of LLMs, we incorporate the SimpleRL dataset [zeng2025simplerl] with 8K mathematical problems for vanilla RL training. For hallucination evaluation, we adopt three widely-used benchmarks, i.e., TruthfulQA [truthfulqa], HaluEval [halueval], and HalluQA [halluqa]. For HaluEval, we adopt the QA subset and report the accuracy with which the model can discriminate correct and hallucinated answers. For TruthfulQA and HalluQA, we report the ratio of outputs that the judge model predicts as truthful. We employ GSM8K [cobbe2021gsm8k], MATH-500 [hendrycksmath2021], AIME 2024 [aime2024], and AIME 2025 [aime2025] to evaluate mathematical reasoning performance and report the Pass@1 accuracy results in these datasets. For more details, refer to Appendix C.
Baselines. We employ Qwen2.5-7B-Base/Instruct [yang2024qwen2] and Llama3.1-8B-Instruct [grattafiori2024llama] as backbone models for our approach. We compare our model against three groups of baselines:
$\bullet$ API-based Models include DeepSeek-V3 [deepseek-v3], DeepSeek-R1 [deepseek-r1], OpenAI GPT-4o, and GPT-o1.
$\bullet$ Reasoning Models consist of QwQ-32B [qwq32b], Distill-Qwen-7B/14B/32B, and Distill-Llama-8B [deepseek-r1], which are fine-tuned on the long CoT data (800K samples) synthesized from DeepSeek-R1.
@@
Implementation Details. We train our model using the verl [sheng2024hybridflow] framework. Specifically, during training, we use a batch size of 8, generate 8 rollouts for each prompt, set a maximum prompt and response length of 2,048 tokens, and train using a mini-batch size of 1,024. We train our model on the mixture dataset of HotpotQA subset and SimpleRL for one epoch with a constant learning rate of 4e-7 and temperature parameter of 1.0. The coefficient for KL loss is set to 1e-3 and the clip ratio is 0.2. We adopt the HHEM-2.1 model [HughesBae2023] to verify step-wise factuality. During evaluation, we set the sampling temperature to 1.0. We present prompt templates throughout this work in Appendix F.
@@ 6.2 Main Results
The experimental results presented in Table 1 demonstrate the performance of our proposed method, FSPO, across various reasoning and hallucination benchmarks compared to multiple competitive models, including API-based, reasoning, and open-source alternatives. On hallucination benchmarks, FSPO clearly outperforms all open-source, reasoning and even some API-based models. Specifically, FSPO based on Qwen2.5-7B-Base achieves the highest scores among all open-source and reasoning models on hallucination datasets, clearly demonstrating the effectiveness of our training approach in substantially reducing hallucinations and enhancing factual correctness.
@@ 6.3 Further Analysis
RL Algorithm Analysis. We further adapt FSPO to the Reinforce++ algorithm [hu2025reinforce++] for validating the adaptability and effectiveness of FSPO beyond GRPO. As shown in Figure 5, FSPO with Reinforce++ exhibits significant improvement over Reinforce++ alone, closely following the performance of FSPO with GRPO. Moreover, FSPO with GRPO consistently achieves higher accuracy than the Reinforce++ baselines across training steps on both the MATH-500 and HaluEval-QA benchmarks. These findings indicate that our proposed FSPO approach generalizes well to other reinforcement learning algorithms, enhancing their effectiveness and stability in solving complex reasoning and reducing hallucinations.
@@
Number of Training Samples. To investigate the model performance under varying numbers of training samples, we randomly sample 1K, 2K, 4K, and 8K samples from the HotpotQA subset [r1-searcher], with the results shown in Figure 7. On MATH-500, the accuracies are similar from 1K to 4K samples while diminishing with 8K samples, which might be because more factual QA data lower the math reasoning performance. In contrast, on the HaluEval-QA task, significant accuracy improvement continues up to 2K samples, with 4K and 8K samples achieving mediocre results. This demonstrates that our approach can significantly reduce hallucinations using only a small amount of data, without compromising the model’s capacity for complex reasoning.